OpenStudy (kirbz):

3/4 = x/28. What is x?

alones (alones):

Step 1: Cross-multiply. 3/4=x/28 (3)*(28)=x*(4) 84=4x Step 2: Flip the equation. 4x=84 Step 3: Divide both sides by 4. 4x/4=84/4 x=21 Btw dear next time post your quesiton math section xP

imqwerty (imqwerty):

welcome to Openstudy!! :D please ask math related question in math section here is the link to math section- http://openstudy.com/study#/groups/mathematics

Hayhayz (hayhayz):

\(\large\color{teal}{Welcome~to~OpenStudy!}\)